Given the vectors A = 8i + 4j -2k N B = 2j + 6k m C = 3i -2j + 4k m Calculate the following: a. 𝐀 βˆ™ 𝐁 b. The orthogonal component of B in the direction of C c. The angle between A and C d. A x B e. a unit vector Ξ» that is perpendicular to both A and B f. 𝐀 x 𝐁 βˆ™ 𝐂


1
Expert's answer
2021-06-01T10:51:04-0400

A = 8i + 4j -2k N

B = 2j + 6k m

C = 3i -2j + 4k m

(a) "A.B= (8i + 4j -2k).(2j + 6k)" Β 

"=(8\times0)+(4\times2)+(-2\times6)"

"=-4"


(b) Orthogonal component of B to C is given by,

"=\dfrac{\vec C\times(\vec B\times\vec C)}{|\vec C|^2}"

"=\dfrac{(3i -2j + 4k)\times((2j + 6k)\times(3i -2j + 4k))}{|3i -2j + 4k|^2}"

"=\dfrac{(3i -2j + 4k)\times(20i+18j-6k)}{29}"

"=\dfrac{-60i+98j+94k}{29}"


(c) Angle between A and C is given by,

"\theta=\cos^{-1}\dfrac{\vec A.\vec C}{|\vec A||\vec C|}"

"\theta=\cos^{-1}\dfrac{(8i + 4j -2k).(3i -2j + 4k)}{|8i + 4j -2k||3i -2j + 4k|}"

"=\cos^{-1}\dfrac{8}{\sqrt{84\times29}}"

"\theta=80.67\degree"


(d) "\vec A\times \vec B=(8i + 4j -2k)\times( 2j + 6k)"

"=28i-48j+16k"


(e) Unit vector perpendicular to both A and B is given by,

"\lambda=\dfrac{\vec A\times\vec B}{|\vec A\times\vec B|}"

"\lambda=\dfrac{28i-48j+16k}{|28i-48j+16k|}"

"\lambda=\dfrac{28i-48j+16k}{3344}"


(f) "(\vec A\times\vec B).\vec C=(28i-48j+16k).(3i -2j + 4k)"

"=244"